# Distribution of local signs of modular forms and murmurations of Fourier coefficients

### v2

@@ 6.1. Analysis for Type I
We want to consider the limit of these averages as $\ell,X\to\infty$ such that $\frac{\ell}{X}\to x$ for some $x\in[0,\infty)$ by substituting the trace formula from 5.5 into (6.1). Specifically, Conjecture 1.8 asserts that the limit exists. Note that one only gets $s$-terms appearing for $s^{2}\leq\frac{4\ell}{Q}\leq\frac{4\ell M}{X}\sim 4Mx$. Consequently, we will only see a bounded number of $s$-terms. As #$\#\mathcal{F}(X,\beta X)$ grows like a multiple of $X^{2}$ (see [zubrilina, Section 3.4] for a precise estimate), the $\delta_{N=1}$ term will contribute nothing asymptotically, and it is easy to see that the $\delta_{k=2}$ term will asymptotically contribute a constant as $\frac{\ell}{X}\to x$.
@@
For a discriminant $-D<-4$, we have $h(-D)=\frac{\sqrt{D}}{\pi}L(1,\chi_{-D})$, where $\chi_{-D}(n)={-D\overwithdelims( )n}$. We have $L(1,\chi_{-D})=\sum_{n<T}\frac{\chi_{-D}(n)}{n}+O(\sqrt{D}\log(D)/T)$ by Polya–Vinogradov and partial summation. Then
@@ 6.2. Analysis for Type II
$-\frac{1}{2}U_{k-2}(\frac{s}{2}\sqrt{\frac{Q}{\ell}})H(\Delta)\sideset{}{{}^{\prime}}{\sum}_{X<M<\beta X}\xi_{\Delta}(M),$
@@ 6.4. Levels divisible by $\ell$
±$\sum_{\mathcal{F}^{\pm}(X,\beta X)}\ell^{1-k/2}a_{\ell}(f)=\sum_{\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}}\ell^{1-k/2}a_{\ell}(f)+O(\frac{X}{\ell}).$
Assuming $\frac{\ell}{X}\to x$, the error term in this expression is $O(1)$ and will go to 00 upon dividing by #±$\#\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}$ or #±$\#\mathcal{F}^{\pm}(X,\beta X)$. Since #±$\#\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}\approx(1-\frac{1}{\ell})\#\mathcal{F}^{\pm}(X,\beta X)$, we see there is no asymptotic difference between working with averages over ±$\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}$ or ±$\mathcal{F}^{\pm}(X,\beta X)$.
@@ 7.1. Twisting at $q$
If $\pi_{q}$ is a ramified supercuspidal (so $r$ is odd) induced from $E_{q}/\mathbb{Q}_{q}$, then $\kappa(\pi_{q},\chi_{q})=\pm 1)$, where the sign is $+1$ if $E_{q}=\mathbb{Q}_{q}(\sqrt{q^{*}})$ and $-1$ if $E_{p}=\mathbb{Q}_{p}(\sqrt{-q^{*}})$.
@@ 7.2. Twisting away from $q$
    there exists an odd $p$ such that $p^{3}\mid N$ and (${q\overwithdelims( )p}=-1$;
@@
Appendix A Errata for “Rank bias for elliptic curves mod $p$” by Kimball Martin and Thomas Pharis

### v3

@@ 6.1. Analysis for Type I
We want to consider the limit of these averages as $\ell,X\to\infty$ such that $\frac{\ell}{X}\to x$ for some $x\in[0,\infty)$ by substituting the trace formula from 5.5 into (6.1). Specifically, Conjecture 1.8 asserts that the limit exists. Note that one only gets $s$-terms appearing for $s^{2}\leq\frac{4\ell}{Q}\leq\frac{4\ell M}{X}\sim 4Mx$. Consequently, we will only see a bounded number of $s$-terms. As $\mathcal{F}(X,\beta X)$ grows like a multiple of $X^{2}$ (see [zubrilina, Section 3.4] for a precise estimate), the $\delta_{N=1}$ term will contribute nothing asymptotically, and it is easy to see that the $\delta_{k=2}$ term will asymptotically contribute a constant as $\frac{\ell}{X}\to x$.
@@
For a discriminant D<4-D<-4, we have $h(-D)=\frac{\sqrt{D}}{\pi}L(1,\chi_{-D})$, where $\chi_{-D}(n)={-D\overwithdelims( )n}$. We have $L(1,\chi_{-D})=\sum_{n<T}\frac{\chi_{-D}(n)}{n}+O(\sqrt{D}\log(D)/T)$ by Polya–Vinogradov and partial summation. Then
@@ 6.2. Analysis for Type II
$\frac{1}{2}U_{k-2}(\frac{s}{2}\sqrt{\frac{Q}{\ell}})H(\Delta)\sideset{}{{}^{\prime}}{\sum}_{X<M<\beta X}\xi_{\Delta}(M),$
@@ 6.4. Levels divisible by $\ell$
$\sum_{\mathcal{F}^{\pm}(X,\beta X)}\ell^{1-k/2}a_{\ell}(f)=\sum_{\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}}\ell^{1-k/2}a_{\ell}(f)+O(\frac{X}{\ell}).$
Assuming $\frac{\ell}{X}\to x$, the error term in this expression is $O(1)$ and will go to 0 upon dividing by $\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}$ or $\mathcal{F}^{\pm}(X,\beta X)$. Since $\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}\approx(1-\frac{1}{\ell})\#\mathcal{F}^{\pm}(X,\beta X)$, we see there is no asymptotic difference between working with averages over $\mathcal{F}^{\pm}(X,\beta X)^{(\ell)}$ or $\mathcal{F}^{\pm}(X,\beta X)$.
@@ 7.1. Twisting at $q$
If $\pi_{q}$ is a ramified supercuspidal (so $r$ is odd) induced from $E_{q}/\mathbb{Q}_{q}$, then $\kappa(\pi_{q},\chi_{q})=\pm 1)$, where the sign is +1+1 if $E_{q}=\mathbb{Q}_{q}(\sqrt{q^{*}})$ and 1-1 if $E_{p}=\mathbb{Q}_{p}(\sqrt{-q^{*}})$.
@@ 7.2. Twisting away from $q$
    there exists an odd $p$ such that $p^{3}\mid N$ and ${q\overwithdelims( )p}=-1$;
@@
Appendix A Errata for “Rank bias for elliptic curves mod pp” by Kimball Martin and Thomas Pharis